A company currently retains 40% of its earnings, which amounts to R6 per share this year. It earns an ROE of 30%. Assuming a required rate of return of 16%, how much would you pay for shares in the company on the basis of the earning multiplier model?
Answer #1

EPS = 6/.40 = R15 per Share since R6 is 40 % of EPS.

Earning Multiplier = 1/Ke = 1/.16= 6.25 times

So as per earning multiplier model=

MPS = EPS*Earning Multiplier

R15*6.25 = R93.75 Per Share